OpenStudy (anonymous):

The angular velocity of a 755-g wheel 15.0 cm in diameter is given by the equation ω(t)=(2.00rad/s2)t+(1.00rad/s4)t3. (a) Through how many radians does the wheel turn during the first 2.00 s of its motion?

OpenStudy (loser66):

\[\theta = \int_0^2 \omega (t) = t^2+\frac{t^4}{4} (from ~~0 ~~to ~~2)\]= 8radians